OpenStudy (anonymous):

Explain how to write the equation of a line if you know its y-intercept and the equation of a line to which it is parallel.

OpenStudy (anonymous):

Because the format is y=mx+b, b=y intercept, u basically copy the mx part of the parrel equation if u know the y intercept

OpenStudy (anonymous):

Just look at this graph, these equations are the same except for the y intercept is different, u just need to copy the mx part and add the intercept to make the equation

OpenStudy (anonymous):

Parrel equations: y=?x+5 y=2/3x+7 the ? would be the 2/3, because they are aprrel and what makes htem parrel is that they have the same mx part and they have diff intercepts

OpenStudy (anonymous):

thanks